OpenStudy (anonymous):

HELP!!! What is the asymptote for the above function y=2x^2/3x^2-16???

OpenStudy (anonymous):

the asymptote is at 3x^2=16, therefore it is sqrt(5&1/3)

OpenStudy (chriss):

the horizontal asymptote is\[\frac{2}{3}\] as x goes to \[\pm \infty\] the vertical asymptote is \[\pm \infty\] as x goes to \[\pm \frac{4}{[\sqrt{3}]}\]